# Grasping a Handful: Sequential Multi-Object Dexterous Grasp Generation

### v4

@@ VI Conditional Sequential Grasp Diffuser
Finally, we introduce SeqDiffuser, a diffusion-based sequential grasp generation method trained on SeqDataset. The architecture of SeqDiffuser is similar to the sampler from [3], with the distinction that SeqDiffuser is also conditioned on the OS to enable sequential multi-object grasping.
@@
where $\hat{\mu}_{\bm{\psi}}(\mathbf{\tilde{g}}_{t},\mathbf{h}_{\mathbf{O}},t)$ and $\hat{\Sigma}_{{\bm{\psi}}}(\mathbf{\tilde{g}}_{t},\mathbf{h}_{\mathbf{O}},t)$ are the learnable mean and variance of a Gaussian distribution and $\mathbf{h}_{\mathbf{O}}$ is the Basis Point Set (BPS[35] encoded feature of the target object $\mathbf{O}$. In the forward and reverse processes, noise is only added or removed at positions where $\mathbf{k}$ is nonzero. The loss function is thereby formulated as:
@@ VII Experiments
We compare our method to the optimization-based sampler MultiGrasp from [13], which generates simultaneous grasps on clustered objects. As such, for MultiGrasp, we must first sample clustered object configurations and then generate multi-object grasps directly on the object cluster. In contrast, SeqGrasp and SeqDiffuser do not require objects to be spatially close as they generate grasps sequentially based on previously successful ones. While this comparison is not entirely fair, we believe that comparing these two strategies still offers valuable insights.
@@ VII-A Simulation Experiments
We evaluated all generated grasps in the simulation experiments in Isaac Gym[36]. We used two object sets: (1) all eight objects from Grasp’Em [13] and (2) a random selection of eight validation objects from SeqDataset. We randomly generated ten four-object sequences for each object set, and, per object, we generated 256 grasps, resulting in 10,240 grasps per method.
@@
We observe a notable decline in performance across all methods when transitioning from three-object to four-object grasps. We hypothesize that this decline occurs because the three previously grasped objects occupy substantial space within the Allegro Hand, pushing the fourth object grasp to the limits of the hand’s kinematic redundancy. Additionally, as the number of grasped objects increases, object-object interactions grow exponentially, making the task considerably more challenging, a finding also reported in [13]. Nevertheless, SeqGrasp demonstrates superior performance in scenarios involving three or more objects.
Another notable observation is that SeqDiffuser generates grasps with high penetration, which aligns with previous work on single-object diffusion-based grasp sampling [3]. Still, SeqDiffuser is valuable because it generates grasps 750-1250 times faster than SeqGrasp and MultiGrasp.
@@ VII-B Real-World Experiments
We followed the procedure outlined in [25] to replicate the grasp on the real hardware. This procedure involves positioning each object as closely as possible with its generated pose and then closing the relevant joints of the hand to retain the objects. For grasps with severe penetration issues, we placed the object in the closest feasible real-world position. For MultiGrasp and SeqGrasp, we replicated the successful grasps with the lowest energy, while for SeqDiffuser, we randomly chose one of the generated grasps. Due to the larger real-world objects, four-object grasping exceeded the workspace limits of the Allegro Hand. Thus, we only evaluated three-object grasps on that object set. In contrast, both three- and four-object grasping were tested on the 3D-printed object set.

### v5

@@ VI Conditional Sequential Grasp Diffuser
Finally, we introduce SeqDiffuser, a diffusion-based sequential grasp generation method trained on SeqDataset. The architecture of SeqDiffuser is similar to the sampler from [dexdiffuser], with the distinction that SeqDiffuser is also conditioned on the OS to enable sequential multi-object grasping.
@@
where $\hat{\mu}_{\bm{\psi}}(\mathbf{\tilde{g}}_{t},\mathbf{h}_{\mathbf{O}},t)$ and $\hat{\Sigma}_{{\bm{\psi}}}(\mathbf{\tilde{g}}_{t},\mathbf{h}_{\mathbf{O}},t)$ are the learnable mean and variance of a Gaussian distribution and $\mathbf{h}_{\mathbf{O}}$ is the Basis Point Set (BPS[prokudin2019efficient] encoded feature of the target object $\mathbf{O}$. In the forward and reverse processes, noise is only added or removed at positions where $\mathbf{k}$ is nonzero. The loss function is thereby formulated as:
@@ VII Experiments
We compare our method to the optimization-based sampler MultiGrasp from [multigrasp], which generates simultaneous grasps on clustered objects. As such, for MultiGrasp, we must first sample clustered object configurations and then generate multi-object grasps directly on the object cluster. In contrast, SeqGrasp and SeqDiffuser do not require objects to be spatially close as they generate grasps sequentially based on previously successful ones. While this comparison is not entirely fair, we believe that comparing these two strategies still offers valuable insights.
@@ VII-A Simulation Experiments
We evaluated all generated grasps in the simulation experiments in Isaac Gym[makoviychuk2021isaac]. We used two object sets: (1) all eight objects from Grasp’Em [multigrasp] and (2) a random selection of eight validation objects from SeqDataset. We randomly generated ten four-object sequences for each object set, and, per object, we generated 256 grasps, resulting in 10,240 grasps per method.
@@
We observe a notable decline in performance across all methods when transitioning from three-object to four-object grasps. We hypothesize that this decline occurs because the three previously grasped objects occupy substantial space within the Allegro Hand, pushing the fourth object grasp to the limits of the hand’s kinematic redundancy. Additionally, as the number of grasped objects increases, object-object interactions grow exponentially, making the task considerably more challenging, a finding also reported in [multigrasp]. Nevertheless, SeqGrasp demonstrates superior performance in scenarios involving three or more objects.
Another notable observation is that SeqDiffuser generates grasps with high penetration, which aligns with previous work on single-object diffusion-based grasp sampling [dexdiffuser]. Still, SeqDiffuser is valuable because it generates grasps 750-1250 times faster than SeqGrasp and MultiGrasp.
@@ VII-B Real-World Experiments
We followed the procedure outlined in [yao2023exploiting] to replicate the grasp on the real hardware. This procedure involves positioning each object as closely as possible with its generated pose and then closing the relevant joints of the hand to retain the objects, as exemplified in Fig. 8. For grasps with severe penetration issues, we placed the object in the closest feasible real-world position. For MultiGrasp and SeqGrasp, we replicated the successful grasps with the lowest energy, while for SeqDiffuser, we randomly chose one of the generated grasps. Due to the larger real-world objects, four-object grasping exceeded the workspace limits of the Allegro Hand. Thus, we only evaluated three-object grasps on that object set. In contrast, both three- and four-object grasping were tested on the 3D-printed object set.